OpenStudy (anonymous):

Find the derivative. f(x) = 1/x^16 Do you use the quotient rule? I'm getting the wrong answer and I'm not sure how to go about doing this.

OpenStudy (anonymous):

no no not the quotient rule if you have a number in the numerator

OpenStudy (anonymous):

write as \(x^{-16}\) and use the power rule

OpenStudy (anonymous):

xn^n-1 x^-16 n= -16

OpenStudy (anonymous):

**nX^n-1

OpenStudy (anonymous):

Arg, I guess I got it wrong still. I didn't realize you had to take it further after the nX^n-1 The solution, for anyone interested: d/dx(1/x^16) = d/dx(x^-16) = -16x^-16-1 = -16x^-17 =-16/x^17